2x^3 + x^2 - 2x - 1
factor out completely.
.

Found 2 solutions by palanisamy, anicole:
Answer by palanisamy(496)

2x^3 + x^2 - 2x - 1
=2x^3 -2x + x^2 - 1
=2x(x^2-1)+1(x^2-1)
=(x^2-1)(2x+1)
=(x+1)(x-1)(2x+1)

Answer by anicole(2)

2x^3 + x^2 - 2x - 1
=2x^3 -2x + x^2 - 1
=2x(x^2-1)+1(x^2-1)
=(x^2-1)(2x+1)
=(x+1)(x-1)(2x+1)